# An Autonomous Vision-Based Algorithm for Interplanetary Navigation

### v2

@@ Abstract
The surge of deep-space probes makes it unsustainable to navigate them with standard radiometric tracking. Autonomous interplanetary satellites represent a solution to this problem. In this work, a vision-based navigation algorithm is built by combining an orbit determination method with an image processing pipeline suitable for interplanetary transfers of autonomous platforms. To increase the computational efficiency of the algorithm, an extended Kalman filter is selected as state estimator, fed by the positions of the planets extracted from deep-space images. An enhancement of the estimation accuracy is performed by applying an optimal strategy to select the best pair of planets to track. Moreover, a novel analytical measurement model for deep-space navigation is developed providing a first-order approximation of the light-aberration and light-time effects. Algorithm performance is tested on a high-fidelity, Earth–Mars interplanetary transfer, showing the algorithm applicability for deep-space navigation.
@@ 1 Introduction
As a new era of deep-space exploration and exploitation is rapidly approaching, the adoption of efficient and sustainable navigation methods becomes increasingly crucial. Traditional ground-based radiometric tracking, while accurate and reliable, heavily depends on limited resources, such as ground stations and flight dynamics teams. This approach is not applicable in the long term to all the spacecraft launched in deep space. There is then an urgent need to enhance the level of navigation autonomy for future interplanetary missions.
@@
Different alternatives enable autonomous navigation capabilities. Among them, the primary ones are autonomous X-ray pulsar-based navigation [1, 2], semi-autonomous radio-based navigation [3], and autonomous vision-based navigation (VBN) [4, 5]. Among these, X-ray navigation requires large detectors and long integration times [6]. One-way radiometric tracking still relies on Earth-based infrastructure. Whereas, VBN is an economical and fully ground-independent solution: it enables determining the probe position by observing the movement of celestial bodies on optical images [6]. VBN stands out as it is also compatible with all mission phases toward celestial bodies: cruise [7, 8, 9, 4], mid-range [10, 11, 12, 13], and close proximity [14], including landing [5, 15]. Consequently, numerous VBN solutions for approach and close proximity have already been developed for various missions. However, VBN algorithms tailored for interplanetary navigation have only undergone onboard testing, with no direct application in probe operations, as evidenced by the validation conducted during the Deep-Space 1 (DS1) mission in 1998 [16]. Nevertheless, the increasing interest in deep-space missions [17], particularly concerning CubeSats [18], is expected to push the adoption of autonomous optical navigation even in far-range scenarios alongside the research in this field. Current research primarily focuses on implementing orbit determination (OD) algorithms to determine the probe state [19, 8, 9, 20, 21]. In Karimi and Mortari [19], innovative angles-only Initial Orbit Determination algorithms are developed, whose output is then used within an extended Kalman filter (EKF) embedding light-effects corrections on the planet position in the measurement model. In Franzese et al. [8], the feasibility of the M-ARGO autonomous deep-space navigation experiment is presented over a low thrust trajectory without considering light time correction. In Andreis et al. [9], an OD algorithm suited to be deployed on a miniaturized processor is developed by studying the most promising EKF implementations for onboard applications. Although these works elaborate on autonomy, there is a gap in the literature regarding the development of a fully integrated pipeline embedding an image-processing procedure for extracting information from deep-space images. In Bhaskaran [22], an IP technique to retrieve beacon information is qualitatively mentioned yet not implemented in a fully integrated simulation, and the effect of the measurement errors on the state estimation is not quantified through simulations. Bhaskaran et al. [7] details the procedure adopted to process the deep-space images of DS1. Due to the long exposure time and high-speed slew rate of the mission, complex image patterns were produced for the point sources. Thus, to retrieve accurately the centroids of the bright objects and the beacon position in the image, computationally heavy multiple cross-correlations were applied, following the approach used for the Galileo mission [23]. In this work, an alternative and computationally lighter approach has been preferred based only on geometrical evaluations following the assumptions of having slower slew rates.
@@
To address these literature gaps, this paper develops an autonomous VBN algorithm intended for use during a deep-space transfer, where the estimation accuracy is improved by applying light-effect corrections and an optimal strategy to select the best pair of beacons to track. The contribution to the state-of-the-art is threefold. First, the extended Kalman filter adopted as OD algorithm [9] is integrated with an IP pipeline suited to deep-space navigation [24]. The work in Andreis et al. [9] and Casini et al. [20] is extended by considering synthetic deep-space images as input. In this way, the measurements are the outcome of an IP procedure rather than being simulated with a mere behavioral model. Consequently, they introduce a more authentic representation of the actual measurement error associated with the image processing and real cameras, resulting in a more accurate reproduction of the state estimation error. Second, the VBN filter is developed for CubeSat applications, thus, particular attention has been paid to the computation capabilities of the navigation algorithm. Third, a novel analytical measurement model for deep-space navigation providing a first-order approximation of light-time correction on beacon position is presented. The proposed model avoids any calculations on the raw camera measurement, so decoupling the spacecraft prediction from the process noise and prevents onboard optimization as in Andreis et al. [9]. Moreover, light-aberration correction is also applied to stars position, being the attitude determined from deep-space images.
@@
The paper is structured as follows. In Sec. 2 the interplanetary navigation problem is described by paying particular attention to the definition of the optimal beacon selection method and light-effects perturbations relevant in the deep-space environment. Sec. 3 details the IP procedure to extract observations from deep-space images. In Sec. 4, the developed VBN filter to be used during an interplanetary transfer is presented. Here, the dynamics and measurement models are described together with the chosen filtering scheme. Eventually, the performance of the IP pipeline and the VBN filter tested on an interplanetary high-fidelity ballistic trajectory is reported in Sec. 5.
@@ 2.1 Problem Geometry
A probe can determine its location by acquiring information from the observation of celestial bodies through optical sensors. Since celestial objects are unresolved in deep space (i.e., they fall within a single pixel) their line-of-sight (LoS) direction or pixel position is the only available information that can be used to estimate the probe state. When two LoS directions associated with different beacons are obtained simultaneously, the kinematic celestial triangulation problem can be solved [19, 25, 4].
@@
    only one miniaturized optical sensor (e.g., star tracker or camera) is adopted;
@@
However, note that the algorithm can be also used for larger spacecraft, despite the examples reported in this paper.
Since to solve the kinematic celestial triangulation problem at least two different synchronous observations are needed, the static celestial triangulation cannot be exploited for the CubeSat operational scenario, given the low probability of detecting several planets with only one instrument. It is worth noting that, in case of a fortunate detection of two planets, the position triangulation would be affected by high uncertainty, given that the two associate LoSes are almost parallel [27]. Therefore, dynamic estimators (e.g., Kalman filtering) are adopted as they can process asynchronous observations and estimate the probe velocity.
@@ 2.2 Optimal Planets Selection
To reach the highest accuracy possible in the state estimation, the approach described in Franzese and Topputo [27] is adopted to optimally select the planets to observe during the interplanetary transfer. The optimal planets pair is chosen among the observable ones by minimizing the figure of merit $\mathcal{J}$, which is the trace of the position error covariance matrix when considering perturbed LoS directions. It is defined as follows:
@@
where $\bm{r}_{i}$ and $\bm{r}_{j}$ are the positions of the two planets in the inertial reference frame, respectively. It is convenient to divide $\bm{d}$ by 1 AU to keep $\mathcal{J}$ non-dimensional.
The optimal planet pair is selected taking into account the planets observability, which is preliminarily assessed by evaluating the planet apparent magnitude and Solar Exclusion Angle (SEA). For more information refer to Andreis et al. [9]. In this work, the optimal planet selection procedure is used to choose the two planets to be sequentially targeted in different images, typically containing only one planet at a time.
@@ 2.3 Light-Effects Perturbations
Another important aspect for deep-space navigation is the impact of light effects (i.e., light time and light aberration [19]) on the observations used to estimate the spacecraft state. The light-effect perturbations are caused by two factors. The first one is caused by the large distance between the spacecraft and the beacon involved in deep space and by the finite speed of the light. This implies that the light detected at the camera has been emitted from the source in the past, causing the celestial object to be observed shifted with respect to its position in the instant of detection by the camera. The fastest and further the planet is, the more significant the light-time effect is. For what concerns the second factor, light aberration is a phenomenon where celestial objects display an apparent motion with respect to their true positions relative to the observer velocity. Since the spacecraft velocity is not negligible in deep space, light-aberration effects become important. The location shift of the celestial body depends on the velocity magnitude and direction relative to the LoS of the observed planet.
These two effects shall be corrected in the filter to avoid systematic errors in the estimation of the spacecraft state. Previous works consider these effects by applying corrections only to the planet LoS directions [19, 9]. Instead, in this work, as the probe attitude is derived from deep-space images, it becomes necessary to apply light aberration correction to the calculated stars LoS directions, as typically done by star trackers [28, 29]. This correction is essential to prevent the computation of a biased attitude value. It is worth noting that only light aberration requires a correction for stars, as they are considered fixed with respect to the Solar System. The corrections of these two effects are described in the upcoming sections.
@@ 3 Image Processing Pipeline for Deep-Space Vision-Based Navigation
Let =$\mathbb{C}=\{C,\bm{C}_{1},\bm{C}_{2}\}$ be the 2D camera reference frame where $\bm{C}_{1}$ points to the right, $\bm{C}_{2}$ downward, and the reference frame center $C$ is placed at the upper left-hand corner of the image. In deep space, the projection of the planet position in $\mathbb{C}$, called ${}^{\mathbb{C}}\bm{r}_{\rm{pl}}$, or its associated LoS direction, is the only information available to support state estimation. An IP algorithm suited for deep-space navigation is adopted to extract this information from the image. The goal of the IP procedure is to recognize the planet locations in the image among the centroids available. The procedure, whose high-level representation is shown in Fig. 1, goes through three steps: 1) The probe attitude is determined, 2) the light-aberration correction is applied to bright star centroids, and 3) the planets are identified. Note that the first step is to identify the portion of the sky the probe is observing and recognize those bright spots that correspond to non-stellar objects in the image. Although the current implementation foresees the attitude determination from the image, note that the Attitude Determination and Control System can also provide this solution in an operative scenario.
@@ 3.1 Attitude Determination
As first step, the probe determines its attitude. To this aim, Niblack’s thresholding method [30] is adopted to remove the background noise to patches of the image centered on bright pixels and delimited by squared windows with a one-pixel margin on each side. Hence, the centroid of the object is computed by applying an intensity-weighted center of gravity algorithm considering the pixels inside the associated squared window [31]. Specifically, the intensity-weighted center-of-gravity method has been selected as it presents enhanced accuracy compared to the simple center-of-gravity approach and minor computation effort compared to Gaussian fitting procedures [32], respecting the computational constraint enforced by the CubeSat application.
Then, the registration problem is solved to find the correct matching between the observed star asterism (i.e., stellar pattern) and the cataloged stars in the inertial frame. This last step is performed differently according to whether the planet is acquired for the first time or not.
3.1.1 Lost-In-Space Strategy
@@
In the former case, the selected Lost-In-Space (LIS) strategy is the pyramid algorithm detailed in Mortari et al. [33]. In this work, the pyramid algorithm has been preferred over the binary search technique [34] for its higher speed gain rate (from 10 to more than 50 times [35]) and for its robustness to spikes (i.e., objects not recognized as stars by the pyramid algorithm). During the operational phase, the pyramid algorithm accesses the star catalog by exploiting an additional vector of integer stored onboard (i.e., the k-vector) which avoids binary search and makes searching time independent from the database size. Specifically, the interstellar angle is the invariant chosen to build the vector adopted for onboard star identification. To meet the stringent memory constraints of a miniaturized onboard computer, solely angles less than 35 degrees are considered to reduce the catalog size. Additionally, only stars with an apparent magnitude lower than 5.5 are adopted in the generation of the invariant. Moreover, the pyramid feature extraction phase is based on subgraph isomorphism. In other terms, an isomorphic subgraph (in this case a pyramid) corresponding to the measured one has to be found in the onboard database for star identification [36]. Note that the objects identified by the pyramid algorithm as spikes may be non-stellar objects (such as planets, asteroids, and cosmic rays) or stars not recognized due to errors in the centroid extraction. Yet, when a great number of spikes is present in the image, the star asterisms may not be recognized by the algorithm. In this work, to reduce the number of scenarios in which this failure occurs, a heuristic approach is considered. Since faint stars are typically not stored in the onboard catalog and centroids extraction is influenced by the thresholding procedure, the procedure is iterated with an increased intensity threshold if attitude determination fails. This approach has the side effect of reducing the number of bright objects in the image, potentially eliminating some spikes. The iteration continues until observed star asterisms are recognized or fewer than three stars are detected.
3.1.2 Recursive Strategy
@@
When the spacecraft is not in LIS mode, it has a rough estimate of its orientation. Therefore, a recursive registration method can be applied. Indeed, by knowing the previous attitude estimation, the LoS directions in the inertial reference frame $\mathcal{I}$ of the four corners of the image are determined. At this point, a check is performed to identify which stars of the onboard catalog are contained inside the image Field of View (FoV). Thus, their locations in the 2D camera reference frame are predicted, and they are associated with the closest centroids of the bright objects extracted from the image when the distances do not exceed a selected threshold value.
@@
When stars are identified, the probe attitude is determined by solving Wahba’s problem [37] between the stars’ LoS directions in the camera and inertial reference frame exploiting the Singular Value Decomposition (SVD) method [37]. Note the camera reference frame coincides with the body reference frame for the sake of simplicity, while the inertial reference frame is the International Celestial Reference Frame (ICRF). Moreover, the robustness of the solution to Wahba’s problem is increased thanks to the adoption of a RANdom-SAmple Consensus (RANSAC) procedure [38, 39]. The RANSAC algorithm is an outlier rejection method aims to detect the bright objects that have been misidentified by the star identification, which can thus lead to a wrong attitude determination. To detect these outliers, the attitude of the spacecraft is adopted as the mathematical model for the data fitting. The attitude is estimated $n_{R}$-times by selecting randomly every time a group of 3 identified stars. A small set of stars needed for attitude determination is chosen to increase the probability of having a group made of different stars at each time. Thus, the estimated $n_{R}$ spacecraft orientations are compared to identify the best model, which is then adopted for the data fitting. The stars not respecting the best model are considered outliers and are labeled as spikes. When the recursive attitude determination fails, the spacecraft orientation at the following image acquisition is determined again with the LIS method. Conversely, when the LIS algorithm succeeds in the determination of the probe orientation, in the following image acquisition the recursive attitude determination algorithm is adopted.
Figure 1: High-level workflow of the two modes of the image processing procedure

### v3

# Please cite this paper as:

@@ Abstract
The surge of deep-space probes makes it unsustainable to navigate them with standard radiometric tracking. Self-driving interplanetary satellites represent a solution to this problem. In this work, a full vision-based navigation algorithm is built by combining an orbit determination method with an image processing pipeline suitable for interplanetary transfers of autonomous platforms. To increase the computational efficiency of the algorithm, a non-dimensional extended Kalman filter is selected as state estimator, fed by the positions of the planets extracted from deep-space images. An enhancement of the estimation accuracy is performed by applying an optimal strategy to select the best pair of planets to track. Moreover, a novel analytical measurement model for deep-space navigation is developed providing a first-order approximation of the light-aberration and light-time effects. Algorithm performance is tested on a high-fidelity, Earth–Mars interplanetary transfer, showing the algorithm applicability for deep-space navigation.
@@ 1 Introduction
As a new era of deep-space exploration and exploitation is rapidly approaching, the adoption of efficient and sustainable navigation methods becomes increasingly crucial. Traditional ground-based radiometric tracking, while accurate and reliable, heavily depends on limited resources, such as ground stations and flight dynamics teams. This approach is unsustainable in the long term. There is then an urgent need to enhance the level of navigation autonomy for future interplanetary missions.
Different alternatives grant autonomous navigation capabilities: autonomous X-ray pulsar-based navigation [1, 2], semi-autonomous radio-based navigation [3], and autonomous vision-based navigation (VBN) [4, 5]. Among these, X-ray navigation requires large detectors and long integration times [6]. One-way radiometric tracking still relies on Earth-based infrastructure. Whereas, VBN is an economical and fully ground-independent solution: it enables determining the probe position by observing the movement of celestial bodies on optical images [6]. In addition, VBN is an approach compatible with all mission phases toward celestial bodies: cruise [7, 8, 9, 4], mid-range [10, 11, 12, 13], and close proximity [14], including landing [5, 15]. Several VBN solutions for approach and close proximity have been already adopted by different missions, e.g., NASA’s OSIRIS-REx [16]. Instead, VBN algorithms for interplanetary navigation have only gone through onboard testing without being directly used for probe operations. An example is the validation performed within the Deep-Space 1 (DS1) mission in 1998 [17]. Nevertheless, in recent years, there has been a growing interest in VBN solutions for deep-space exploration, applied in particular to CubeSats missions [6]. Worth mentioning is the Miniaturised Asteroid Remote Geophysical Observer (M-ARGO), which aims to execute an onboard autonomous navigation test during the interplanetary transfer [18].
@@
In interplanetary VBN, previous research primarily focuses on implementing orbit determination (OD) algorithms to determine the probe state [19, 8, 9, 20, 21]. In [19], innovative angles-only Initial Orbit Determination algorithms are developed, whose output is then used within an extended Kalman filter (EKF) embedding light-effects corrections on the planet position in the measurement model. In [8], the feasibility of the M-ARGO autonomous deep-space navigation experiment is presented. In [9], an OD algorithm suited to be deployed on a miniaturized processor is developed by studying the most promising EKF implementations for onboard applications. Although the above works elaborate on autonomous OD, there is less literature focusing on developing a fully integrated pipeline embedding an image-processing (IP) procedure for extracting information from deep-space images. In [22], an IP technique to retrieve beacon information is qualitatively mentioned yet not implemented in a fully integrated simulation, and the effect of the measurement errors on the state estimation is not quantified through simulations. Whereas, [7] details the procedure adopted to process the deep-space images of DS1. Due to the long exposure time and high-speed slew rate of the mission, complex image patterns were produced for the point sources. Thus, to retrieve accurately the centroids of the bright objects and the beacon position in the image, computationally heavy multiple cross-correlations were applied, following the approach used for the Galileo mission [23]. In this work, an alternative and computationally lighter approach has been preferred based only on geometrical evaluations following the assumptions of having slower slew rates.
@@
This work develops an autonomous VBN algorithm intended for use during a deep-space transfer, where the estimation accuracy is improved by applying light-effect corrections and an optimal strategy to select the best pair of beacons to track. The contribution to the state-of-the-art is threefold. First, the extended Kalman filter adopted as OD algorithm [9] is integrated with an IP pipeline suited to deep-space navigation [24]. The literature in [9, 20] is extended by considering deep-space images as input. In this way, the measurements are the outcome of an IP procedure rather than a mere behavioral model, which yields a more realistic representation of the application case and a faithful reproduction of the state estimation error. Second, the VBN filter is developed for CubeSat applications, thus, particular attention has been paid to the computation capabilities of the navigation algorithm. Third, a novel analytical measurement model for deep-space navigation providing a first-order approximation of light-effects correction on beacon position is presented. The proposed model avoids correcting the raw camera measurement, so decoupling the spacecraft prediction from the process noise and prevents onboard optimization as in [9]. Moreover, light-aberration correction is also applied to the stars position, being the attitude determined from deep-space images.
@@
The paper is structured as follows. In Sec. 2 the interplanetary navigation problem is described by paying particular attention to the definition of the optimal beacon selection method and the light-effects perturbations relevant in the deep-space environment. Sec. 3 details the IP procedure to extract observations from deep-space images. In Sec. 4, the developed VBN filter to be used during an interplanetary transfer is presented. Here, the dynamics and measurement models are described together with the chosen filtering scheme. Eventually, the performance of the IP pipeline and of the VBN filter tested on an interplanetary high-fidelity ballistic trajectory is reported in Sec. 5.
@@ 2.1 Problem Geometry
A probe can determine its location by acquiring information from the observation of celestial bodies through optical sensors. Since celestial objects are unresolved in deep space, i.e., they fall within a single pixel, their line-of-sight (LoS) direction or pixel position is the only available information that can be used to estimate the probe state. When two LoS directions associated with different beacons are obtained simultaneously, the kinematic celestial triangulation problem can be solved [19, 25, 4].
@@
    only one miniaturized optical sensor, e.g., star tracker or camera, is adopted;
@@
However, note that the algorithm can be used also for larger spacecraft, despite the application of the paper.
Since the kinematic celestial triangulation problem requires at least two different synchronous observations to be solved, but it is improbable to detect several planets with one single instrument due to their space sparsity, the static celestial triangulation cannot be exploited for the CubeSat operational scenario. Therefore, dynamic estimators, e.g., Kalman filtering, are adopted as they can process asynchronous observations.
@@ 2.2 Optimal Planets Selection
To reach the highest accuracy possible in the state estimation, the approach described in [27] is adopted to optimally select the planets to observe during the interplanetary transfer. The optimal planets pair is chosen among the observable ones by minimizing the figure of merit $\mathcal{J}$, which is the trace of the position error covariance matrix when considering perturbed LoS directions. It is defined as follows:
@@
where $\bm{r}_{i}$ and $\bm{r}_{j}$ are the positions of the two planets, respectively. It is convenient to divide $\bm{d}$ by 1 AU to keep $\mathcal{J}$ non-dimensional.
The optimal planet pair is selected taking into account the planets observability, which is preliminarily assessed by evaluating the planet apparent magnitude and Solar Exclusion Angle (SEA). For more information refer to [9].
@@ 2.3 Light-Effects Perturbations
Another important aspect to consider for deep-space navigation is the impact of light effects, i.e., light time and light aberration [19], on the observations used to estimate the spacecraft state. On one hand, the light-time effect is caused by the fact that, given the large distance involved in deep space and the finite speed of the light, the light detected at the camera is emitted from the target in the past. This brings the celestial object to be observed shifted with respect to its position in the instant of detection. The further the planet is from the spacecraft, the more significant the light-time effect is. On the other hand, the light-aberration effect is caused by the relative motion between the observer and the light source, and it becomes important when the spacecraft velocity is not negligible. Like the light-time effect, this causes a change in the planet position projection, which depends on the velocity intensity and direction relative to the LoS of the observed planet.
These two effects shall be corrected in the filter to avoid systematic errors in the estimation of the spacecraft state. Previous works consider these effects by applying corrections only to the planet LoS directions [19, 9]. Instead, in this work, since also the probe attitude is determined from deep-space images, the light-effect corrections need to be applied also to the computed stars LoS directions to avoid the evaluation of a biased attitude value. However, since stars are assumed to be fixed with respect to the Solar System, only light aberration needs to be corrected.
@@ 3 Image Processing Pipeline for Deep-Space Vision-Based Navigation
In deep space, the projection of the planet position in the 2D camera reference frame $\mathbb{C}$, i.e., ${}^{\mathbb{C}}\bm{r}_{\rm{pl}}$, or its associated LoS direction, is the only information available to support state estimation. An IP algorithm suited for deep-space navigation is adopted to extract this information from the image. The goal of the IP procedure is to recognize the planet projections in the image among the centroids available. The procedure goes through three steps: 1) The probe attitude is determined, 2) the light-aberration correction is applied to bright star centroids, and 3) the planets are identified. Note that the first step is needed to identify the portion of the sky the probe is observing and recognize those bright spots that correspond to not-stellar objects in the image. Although the current implementation foresees the attitude determination from the image, note that the Attitude Determination and Control System can also provide this solution in an operative scenario. A graphical representation of the IP procedure is shown in Fig. 1.
@@ 3.1 Attitude Determination
As the first step, the probe determines its attitude. At this aim, Niblack’s thresholding method [28] is adopted to remove the background noise to portions of the image centered on bright pixels and delimited by squared windows with a margin of one pixel on each side. Hence, the centroid of the object is computed by applying an intensity-weighted center of gravity algorithm considering the pixels inside the associated squared window [29]. At this point, the registration problem, whose goal is to find the correct matching between the observed star asterism and the cataloged stars in the inertial frame, is solved. This last step is performed differently according to whether the planet is acquired for the first time or not.
@@
In the former case, the selected lost-in-space (LIS) strategy is the search-less algorithm (SLA) introduced in [30]. In this work, the SLA has been preferred over the binary search technique [31] for its higher speed gain rate (from 10 to more than 50 times [32]) and for its robustness to spikes. To be adopted, the SLA requires computing on the ground a vector of integers, the k-vector, which contains information for the stars matching starting from the chosen stars invariant. In this work, the interstellar angle is the invariant chosen to build the star catalog. To reduce the size of the catalog, only stars whose apparent magnitude is lower than 5.5 are considered for the generation of the invariant. Moreover, interstar angles greater than 35 deg are not taken into account. The objects identified by the SLA as spikes may be non-stellar objects (such as planets, asteroids, and cosmic rays) or stars not recognized due to errors in the centroid extraction. Yet, when a great number of spikes is present in the image, the star asterisms may not be recognized by the algorithm. In this work, to reduce the number of scenarios in which this failure occurs, a heuristic approach is considered. As faint stars are generally not stored in the onboard catalog and as the centroids extraction depends on the thresholding procedure, when the attitude determination fails, the attitude determination procedure is iterated again by increasing the intensity threshold. One of the results of this approach consists of diminishing the number of bright objects in the image, which can ultimately lead to the removal of some spikes. The procedure is repeated until observed star asterisms are recognized or less than three stars are detected.
@@
When the spacecraft is not in LIS mode, it has a rough estimate of its orientation. Therefore, a recursive registration method can be applied. Indeed, by knowing the previous attitude estimation, the LoS directions in the inertial reference frame $\mathcal{N}$ of the four corners of the image are determined. At this point, a check is performed to identify which stars of the onboard catalog are contained inside the image Field of View (FoV). Thus, their position projections in the 2D camera reference frame are evaluated, and they are associated with the closest centroids of the bright objects extracted from the image.
@@
When stars are identified, the probe attitude is determined by solving Wahba’s problem [33] between the stars’ LoS directions in the camera and inertial reference frame exploiting the Singular Value Decomposition (SVD) method [33]. Moreover, the robustness of the solution to Wahba’s problem is increased thanks to the adoption of a RANdom-SAmple Consensus (RANSAC) procedure [34, 35]. The RANSAC algorithm aims to detect the bright objects that have been misidentified by the star identification, which can thus lead to a wrong attitude determination. To detect these outliers, the attitude of the spacecraft is adopted as the mathematical model for the data fitting. The attitude is estimated $n_{R}$-times by selecting randomly every time a group of 3 identified stars. The minimum set of stars needed for attitude determination is chosen to increase the probability of having a group made of different stars at each time. Thus, the estimated $n_{R}$ spacecraft orientations are compared to identify the best model, which is then adopted for the data fitting. The stars not respecting the best model are considered outliers and are labeled as spikes.
When the recursive attitude determination fails, the spacecraft orientation at the following image acquisition is determined again with the LIS method. Vice versa, when the LIS algorithm succeeds in the determination of the probe orientation, in the following image acquisition the recursive attitude determination algorithm will be adopted.
Figure 1: Image Processing General Workflow